# DUNE: Distilling a Universal Encoder from Heterogeneous 2D and 3D Teachers

### v1

@@ 1 Introduction
To evaluate our framework, we choose three state-of-the-art models. Two are highly task-specific: MASt3R [30] solves 3D scene reconstruction and matching while Multi-HMR [4] solves 3D human perception. As a third teacher, we add DINO-v2 [38], a popular visual foundation model that generalizes to various visual downstream tasks including semantic segmentation or monocular depth estimation. This leads to DUNE, obtained by Distilling a UNiversal Encoder from heterogeneous 2D and 3D Teachers.
@@ 2 Related work
In this paper, we explore distillation from heterogeneous teachers, i.e. trained on distinct domains and solving diverse tasks. Specifically, we distill from MASt3R [30] for binocular 3D tasks, Multi-HMR [4] for human mesh recovery, and DINO-v2 [38] for 2D tasks like monocular depth and semantic segmentation. This results in an encoder that excels across many these tasks out of the box.
@@ 3 Problem definition and challenges
    Their individual training set constitutes heterogeneous data. We want to jointly distill teachers trained on huge generic image datasets crawled from the web, together with teachers trained on highly curated and potentially carefully annotated datasets composed of natural or synthetic images.
@@ 4.3 Co-distillation with heterogeneous data
  • Generic data sharing: Projector $h_{i}$ receives only ${\mathcal{D}}_{i}$, i.e. the data associated with ${\mathcal{T}}_{i}$ as well as generic data ${\mathcal{D}}_{g}$.
@@ 5.1 Evaluation protocol
Table 1: Distillation data and projector design. We distill models using either ImageNet-19K as a generic dataset or all the 19 teacher datasets combined with full data sharing (Sec. 4.3), employing either a Simple Projector (SP) [42], a Ladder of Projectors (LP) [47] or the Transformer Projector (TP) presented in Sec. 4.2.
@@ Acknowledgements.
We thank the 3DH and GeoDe teams at NAVER LABS Europe who created MASt3R and Multi-HMR, respectively, for sharing their dataset, training and evaluation pipelines. We also thank Cesar de Souza for developing a real-time multi-task demo interface for DUNE, as well as for naming our model.
@@ Appendix B Details on decoder fine-tuning
Semantic segmentation and depth estimation are dense prediction tasks, both formulated as classification tasks in this work, and solved following the simple setup proposed in [38], also followed by the most recent related works [42, 47]. We extract the tokens from the last output layer of the student model and use as input to a linear prediction head. For semantic segmentation, we additionally use the Transformer Projector of the DINO-v2 teacher as part of the frozen encoder, and train a linear head on top of the projector to predict class logits from a patch token. This yields a $32{\times}32$ logit map that is upsampled via bilinear interpolation to the original image resolution of $512{\times}512$.
@@ D.1 Multi-view depth evaluation
We follow the protocol of [49] and evaluate multi-depth stereo depth evaluation on KITTI [22], DTU [2], ETH3D [48], Tanks And Temples [29] and ScanNet [13]. We report the Absolute Relative Error (rel) and the Inlier Ratio ($\tau$) with a threshold of $1.03$ on each test set, as well as the averages over all test sets. To extract depth prediction of one image, we follow DUSt3R [62] and extract depthmaps as the z-coordinate of the predicted pointmaps; and when multiple pointmaps are available for one image from different image pairs, we simply rescale the predicted depthmaps and average them with weights given by the predicted confidence values. Results are reported in Table 7. DUNE performs similarly to MASt3R and DUSt3R on this task overall, while using a smaller ViT-Base image encoder. There is a small decrease of performance on KITTI, which might be caused by the few images from driving dataset when distilling the encoder.
Method KITTI ScanNet ETH3D DTU T&T Average
rel. $\downarrow$ $\tau$uparrow rel. $\downarrow$ $\tau$uparrow rel. $\downarrow$ $\tau$uparrow rel. $\downarrow$ $\tau$uparrow rel. $\downarrow$ $\tau$uparrow rel. $\downarrow$ $\tau$uparrow
DeepV2D [56] 10.00 36.20 4.40 54.80 11.80 29.30 7.70 33.00 8.90 46.40 8.60 39.90
DUSt3R [62]   9.11 39.49 4.93 60.20   2.91 76.91 3.52 69.30 3.17 76.68 4.73 64.52
MASt3R [30]   9.62 46.95 5.86 54.96   2.44 82.77 3.46 66.89 2.04 87.88 4.68 67.89
DUNE 11.36 33.83 6.10 48.97   2.48 77.97 2.70 75.63 2.60 79.19 5.05 63.12
@@ D.2 Multi-view camera pose regression evaluation
Method Co3Dv2 $\uparrow$ RealEstate10K $\uparrow$
RRA@15 RTA@15 mAA(30) mAA(30)
DUSt3R [62] 93.3 88.4 77.2 61.2
MASt3R [30] 94.6 91.9 81.8 76.4
DUNE 92.2 90.7 78.8 79.9
@@
Figure 8: Evaluating DUNE on the Feat2GS benchmark. The spider plot shows comparisson of different encoder models. The Feat2GS benchmark [10] evaluates Novel View Synthesis as a proxy for 3D awareness. In all metrics, larger distance to the center indicates better performance.
@@ D.3 Evaluating DUNE on the Feat2GS benchmark
In Fig. 8 we compare different encoder models in the Feat2GS benchmark [10]. The Feat2GS benchmark has three modalities, i) Geometry: When only geometry parameters are predicted from features and texture is free-optimized for Novel View Synthesis. ii) Texture: When only the texture is predicted from encoder features and the geometry is free-optimized. And iii) All: When both geometry and texture are predicted from features. Our DUNE encoder leads to the best performance when All the parameters are predicted from features (to our understanding the most challenging setting) and leads to the largest area over all settings and metrics. For more detailed results, we also present Tab. 11 with per-dataset evaluations of all metrics and modalities.
@@
D.4 Comparison to 3D-to-2D distillation and 3D-uplifting methods
In Tab. 9, we report semantic segmentation evaluations on three datasets, comparing DUNE to Pri3D [24] (a 3D-to-2D distillation method) and MASt3R. For DUNE, we present results using the encoder outputs directly (DUNE) and with the DINO-v2 projector applied after the encoder (DUNE $+$ DINO-v2 projector). In all cases, only a linear layer is trained to predict patch labels. DUNE significantly outperforms both Pri3D and MASt3R.
@@ D.4 Comparison to 3D-to-2D distillation and 3D-uplifting methods
Table 11: Per-dataset results of Novel View Synthesis metrics in the Feat2GS benchmark [10].
@@ D.5 Qualitative comparisons of teacher outputs to DUNE
In Figs. 10, 11 and 12 we present qualitative results for the MASt3R teacher and our student side-by-side. We see that the student clearly improves over the teacher in some cases.

### v2

@@ 1 Introduction
To evaluate our framework, we choose three state-of-the-art models. Two are highly task-specific: MASt3R [30] solves 3D scene reconstruction and matching, while Multi-HMR [4] solves 3D human perception. As a third teacher, we add DINO-v2 [38], a popular visual foundation model that generalizes to various visual downstream tasks including semantic segmentation or monocular depth estimation. This leads to DUNE, obtained by Distilling a UNiversal Encoder from heterogeneous 2D and 3D Teachers.
@@ 2 Related work
In this paper, we explore distillation from heterogeneous teachers, i.e. trained on distinct domains and solving diverse tasks. Specifically, we distill from MASt3R [30] for binocular 3D tasks, Multi-HMR [4] for human mesh recovery, and DINO-v2 [38] for 2D tasks like monocular depth and semantic segmentation. This results in an encoder that excels across many of these tasks out of the box.
@@ 3 Problem definition and challenges
    Their individual training sets consist of heterogeneous data. We want to jointly distill teachers trained on huge generic image datasets crawled from the web, together with teachers trained on highly curated and potentially carefully annotated datasets composed of natural or synthetic images.
@@ 4.3 Co-distillation with heterogeneous data
  • Generic data sharing: Projector $h_{i}$ receives only ${\mathcal{D}}_{i}$ (the data associated with ${\mathcal{T}}_{i}$) and generic data ${\mathcal{D}}_{g}$.
@@ 5.1 Evaluation protocol
Table 1: Distillation data and projector design. We distill models using either ImageNet-19K as a generic dataset or all the 19 teacher datasets combined with full data sharing (Sec. 4.3), employing either a Simple Projector (SP) [42], a Ladder of Projectors (LP) [47], or the Transformer Projector (TP) presented in Sec. 4.2.
@@ Acknowledgements.
We thank the 3DH and GeoDe teams at NAVER LABS Europe who created Multi-HMR and MASt3R, respectively, for sharing their dataset, training and evaluation pipelines. We also thank César de Souza for developing a real-time multi-task demo interface for DUNE, as well as for naming our model.
@@ Appendix B Details on decoder fine-tuning
Semantic segmentation and depth estimation are dense prediction tasks, both formulated as classification tasks in this work, and solved following the simple setup proposed in [38], also followed by the most recent related works [42, 47]. We extract the tokens from the last output layer of the student model and use as input to a linear prediction head. For semantic segmentation, we additionally use the Transformer Projector of the DINO-v2 teacher as part of the frozen encoder, and train a linear head on top of the projector. to predict class logits from a patch token. This yields a $32{\times}32$ logit map that is upsampled via bilinear interpolation to the original image resolution of $512{\times}512$.
@@ D.1 Multi-view depth evaluation
We follow the protocol of [49] and evaluate multi-depth stereo depth evaluation on KITTI [22], DTU [2], ETH3D [48], Tanks And Temples [29] and ScanNet [13]. We report the Absolute Relative Error (rel) and the Inlier Ratio ($\tau$) with a threshold of $1.03$ on each test set, as well as the averages over all test sets. To extract depth prediction of one image, we follow DUSt3R [62] and extract depthmaps as the z-coordinate of the predicted pointmaps; and when multiple pointmaps are available for one image from different image pairs, we simply rescale the predicted depthmaps and average them with weights given by the predicted confidence values. Results are reported in Table 7. DUNE performs similarly to MASt3R and DUSt3R on this task overall, while using a smaller ViT-Base image encoder.
Method Encoder KITTI ScanNet ETH3D DTU T&T Average rel. $\downarrow$ $\tau$uparrow rel. $\downarrow$ $\tau$uparrow rel. $\downarrow$ $\tau$uparrow rel. $\downarrow$ $\tau$uparrow rel. $\downarrow$ $\tau$uparrow rel. $\downarrow$ $\tau$uparrow DeepV2D [56] Hourglass 10.00 36.20 4.40 54.80 11.80 29.30 7.70 33.00 8.90 46.40 8.60 39.90 DUSt3R [62] ViT-Large  5.88 47.67 3.01 72.54 3.04 75.17 2.92 73.94 2.93 78.51 3.56 69.56 MASt3R [30] ViT-Large  3.54 65.68 4.17 65.22 2.44 82.77 3.46 66.89 2.04 87.88 3.13 73.69 DUNE ViT-Base  4.88 50.76 4.24 59.68 2.48 77.97 2.69 75.63 2.60 79.19 3.38 68.65
@@ D.2 Multi-view camera pose regression evaluation
Method Encoder Co3Dv2 $\uparrow$ RealEstate10K $\uparrow$ RRA@15 RTA@15 mAA(30) mAA(30) DUSt3R [62] ViT-Large 93.3 88.4 77.2 61.2 MASt3R [30] ViT-Large 94.6 91.9 81.8 76.4 DUNE ViT-Base 92.2 90.7 78.8 79.9
@@
Figure 8: Evaluating DUNE on the Feat2GS benchmark. The spider plot shows comparisson of different encoder models. The Feat2GS benchmark [10] evaluates Novel View Synthesis as a proxy for 3D awareness. In all metrics, larger distance to the center indicates better performance. Note that models vary in size: RADIOv2 is a ViT-H, MASt3R a ViT-L and DINO-v2 and DUNE a ViT-B.
@@ D.3 Evaluating DUNE on the Feat2GS benchmark
In Fig. 8 we compare different encoder models in the Feat2GS benchmark [10]. The Feat2GS benchmark has three modalities, i) Geometry: When only geometry parameters are predicted from features and texture is free-optimized for Novel View Synthesis. ii) Texture: When only the texture is predicted from encoder features and the geometry is free-optimized. And iii) All: When both geometry and texture are predicted from features. Our DUNE encoder leads to the best performance when All the parameters are predicted from features (to our understanding the most challenging setting) and leads to the largest area over all settings and metrics. For more detailed results, we also present Tab. 11 with per-dataset evaluations of all metrics and modalities. While all encoders use a ViT architecture, they vary significantly in size, mainly due to the absence of ViT-B models for certain methods. Namely, RADIOv2 only has a ViT-H model open-sourced and MASt3R a ViT-L, DINO-v2 and our model DUNE are ViT-B. Thus, the fact DUNE is obtaining the overall best performance compared to much larger models is even more remarkable.
@@
Table 11: Per-dataset results of Novel View Synthesis metrics in the Feat2GS benchmark [10]. Note that models vary in size: RADIOv2 is a ViT-H, MASt3R a ViT-L and DINO-v2 and DUNE a ViT-B.
D.4 Comparison to 3D-to-2D distillation and 3D-uplifting methods
In Tab. 9, we report semantic segmentation evaluations on three datasets, comparing DUNE to Pri3D [24] (a 3D-to-2D distillation method) and MASt3R. For DUNE, we present results using the encoder outputs directly, DUNE (no proj.), and with the DINO-v2 projector applied after the encoder, DUNE. In all cases, only a linear layer is trained to predict patch labels. DUNE significantly outperforms both Pri3D and MASt3R.
@@ D.5 Qualitative comparisons of teacher outputs to DUNE
In Figs. 10, 11 and 12 we present qualitative results for MASt3R and our student side-by-side. We see that the student clearly improves over the teacher in some cases.